How many different wavelengths would appear in the spectrum of hydrogen atoms initially in the n=5 state?



1
Expert's answer
2021-12-03T12:46:11-0500



The number of different transitions from initial state "n=5" to the ground state is given by

"\frac{n(n-1)}{2}=\frac{5(5-1)}{2}=10"